OpenStudy (anonymous):

Elizabeth has worked for 5 months and earned 6 vacation days. If she continues to earn vacation days at this rate, how many more months does she need to work to reach a total of 10 vacation days?

OpenStudy (anonymous):

Help please due tomorrow..

OpenStudy (anonymous):

31/2 I think!

OpenStudy (anonymous):

20/6 I think (I got it from setting up a ratio, 5months/6days and the ratio xmonths/4(more)days. cross mulitply to get 6x=20, then divide both sides to get 20/6)

OpenStudy (anonymous):

I think you would set up a proportion but you would write: 5/6=x/10 cross multiply 6x=50 divide by six 50/6 (8.333)

OpenStudy (anonymous):

Because for 6 days she worked 5 months and for 10 days she worked x months.

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

@iluvlucy that's right, but it says a TOTAL of 10 days, and she already has 6 from her 5 months, so wouldn't you just do the 4 MORE days she needs, instead of all ten? or subtract the 5 months that she already has from your answers. Both would give the answer 20/6 (3.333 months)

OpenStudy (anonymous):

Maybe you're right.... I just remembered this from algebra.... that was years ago.

OpenStudy (anonymous):

Yeah I think you're right